# A Data-Driven Aggressive Autonomous Racing Framework Utilizing Local Trajectory Planning with Velocity Prediction

### v1

@@ V EXPERIMENTAL RESULTS AND DISCUSSION
In this section, the superiority of the proposed VPBO-RF is validated through simulations and real-world experiments. The focus is on BO training efficiency in simulations and the vehicle’s mean velocity, lap time (L.T), and computation time (C.T) during real-world racing using a self-built 1:10 scale F1TENTH vehicle (see Fig. 4a). A complex racetrack with four significant curvature corners (as shown in Fig. 4b) is built for simulation and real-world racing. The Cartographer ROS[23] is used for mapping and localization (see Fig. 4c).
@@
Figure 5: The different colored boxes    represent the optimal lap time obtained from training in Fig. 5a. The proposed Objective Function adapted to Racing enables BO to converge rapidly, and its optimal lap time outperforms the baseline. Ablation experiments demonstrate the effectiveness of different components of OFR for improving the training efficiency and safety of optimal trajectories.
V-A Effectiveness of the proposed OFR
@@
The training efficiency is calculated as $\frac{N_{\textrm{BL}}-N_{\textrm{opt}}}{N_{\textrm{BL}}}$, where $N_{\textrm{opt}}$ is iterations required for BO convergence of different experiments. The percentage of racing performance improvement is calculated as $\frac{T_{\textrm{lap,BL}}-T_{\textrm{lap,opt}}}{T_{\textrm{lap,BL}}}$, where $T_{\textrm{lap,opt}}$ is the optimal lap time. The lower bound of $\bm{\theta}$ is $[5,1,1,1,1,0.1,1,1,0.01]$, and the upper bound is $[60,50,10,10,10,20,50,20,0.4]$. The $D_{\textrm{ref}}$ is 62.8m and ${J}_{\textrm{inf}}$ is 17.6. The results of BO learning based on different objective functions are shown in Fig. 5a and Table II. Compared with baseline, OFR significantly improves the training efficiency of BO by 42.86%. In addition, the OFR converged lap time is also superior to baseline with a reduction of 9.07%. Ablation1 and Ablation2 show lower training efficiency and less reduction in optimal lap time than the proposed OFR. This demonstrates the effectiveness of the trajectory length minimization and lap time reduction components of the proposed OFR. The training efficiency of Ablation3 is comparable to that of the proposed OFR, and it achieves a shorter lap time. However, the slip angle of the optimal trajectory from Ablation3 is significantly larger than that obtained through OFR-based training (as shown in Fig. 5b). This suggests that limiting the maximum orthogonal distance of the racing trajectory can reduce slip angle. This is because OFR penalizes BO for exploring racing trajectories that deviate significantly from the reference line during training.
@@
Methods1 Mean(L.T) Compare2$\downarrow$ Mean($v$) Compare3$\uparrow$ Max($v$) Laps Std(C.T) Mean(C.T) Std(Cost)    Std($a_{x}$)
@@
  • 1 The performance of trajectories[11] with globally minimal curvature, based on ideal vehicle handling capabilities, is regarded as vehicle limits.
@@
  • 3 Calculated as $\frac{{\textrm{Mean}(v_{\textrm{lap}})}}{{\textrm{Mean}(v_{\textrm{limits}})}}$. VPMPCC pushes vehicles to limits through the trade-off between velocity prediction and projected velocity maximization.
(a) Racing trajectories correspond to three different methods.
(b) Longitudinal velocities correspond to racing trajectories.
@@
Figure 6: A schematic of trajectories and vehicle velocities during a single racing lap. The racing trajectory diagram reveals that the locally optimal cornering trajectory differs from the reference line associated with the RVP, as offline trajectory planning cannot account for the real-time states of the vehicle. Furthermore, the velocity diagram shows that VPMPCC achieves the highest cornering velocities, benefiting from its balance between RVP-based velocity prediction and projected velocity maximization in planning optimal local trajectories.
V-B Effectiveness of the proposed VPMPCC
@@
The MPCC proposed in [14] and the Velocity-Tracking Adaptive Pure Pursuit (VTAPP) controller proposed in [20] are used for comparison. The optimal parameters of MPCC are also trained through OFR-based BO, which is $[13,0,39.0,5.8,0.6,2.8,0.2,0.5,0.02]$. And MPCC uses the racetrack centerline as the reference line. The optimal parameters of VPMPCC trained by BO based on the proposed OFR is $\bm{\theta}^{*}=[6,3,6,3.9,1,19.0,28,15.7,0.3]$. These three methods are used to perform consecutive laps of racing to verify the superiority of the VPMPCC method in terms of high performance and stability. The statistical data is shown in Table III.
@@
The mean velocity of VPMPCC reaches 93.18% of vehicle handling limits, much higher than the other methods. It also has a low computation time of 7.04ms, which indicates that VPMPCC has high real-time performance. VPMPCC has consistently raced for 25 laps at a maximum velocity of 6.39 m/s, demonstrating its stability. This successfully proves that VPMPCC pushes the vehicle performance to its limits with stability. Compared to VTAPP, VPMPCC maintains the effect of maximum projected velocity, thus planning local trajectories with superior velocities, especially in corners (as shown in Fig. 6a). As shown in Fig. 6b, during a single lap, the racing velocities of VPMPCC are significantly higher than those of MPCC. This is because MPCC employs an aggressive racing strategy, but the tracking error $\mathbf{e}_{s}$ is not sufficient to enable it to decelerate in time in corner ‘4’, resulting in a possible collision. As a result, the high-velocity racing parameters of MPCC are restricted in BO training, leaving only parameters with limited racing effects. However, VPMPCC can successfully plan optimal local trajectories in corners by incorporating velocity prediction. This reduces cornering time while ensuring collision-free maneuvers, allowing VPMPCC to achieve higher global racing velocities and ultimately reduce lap time. This is also proved by optimal costs of MPCC and VPMPCC in Fig. 7, while MPCC successfully passes corner ‘4’, its parameters cause a premature deceleration in corners ‘1’ to ‘3’ (pointed out by $\uparrow$ in Fig. 6b), leading to a loss in racing performance. In contrast, VPMPCC achieves an optimal velocity profile at each sharp corner. Additionally, the cost changes of VPMPCC are smoother compared to MPCC, which further enhances the efficiency of solving the optimization problem (as shown in Fig. 5c ).
@@ VI CONCLUSIONS
Figure 7: Demonstration of the cost for MPCC and VPMPCC in real vehicle racing. High costs indicate lower velocities. MPCC cannot plan optimal velocity profiles at corners, leading to premature deceleration at corners ‘1’ to ‘3’. In contrast, VPMPCC utilizes velocity prediction to balance projected velocity maximization, maintaining optimal velocities at each corner, thus maximizing racing performance.
This paper proposes a novel data-driven framework for aggressive autonomous racing that pushes the vehicle to its handling limits. By integrating velocity prediction into local trajectory planning, the vehicle’s mean velocity reaches 93.18% of its handling limits. The training efficiency of BO is improved by 42.86% through the proposed objective function adapted to racing. By filtering out unqualified racing trajectories during BO training, the parameters trained in simulation can be directly transferred to a real-world vehicle without retraining. Future work will explore applying VPMPCC to complex multi-vehicle racing scenarios and integrating deep reinforcement learning for decision-making.

### v2

@@ V EXPERIMENTAL RESULTS AND DISCUSSION
Figure 5: The different colored boxes    represent the optimal lap time obtained from training in Fig. 5a. The curves indicate the current best points during training. The proposed Objective Function adapted to Racing enables BO to converge rapidly, and its optimal lap time outperforms the baseline. Ablation experiments demonstrate the effectiveness of different components of OFR for improving the training efficiency and safety of optimal trajectories.
In this section, the superiority of the proposed VPBO-RF is validated through simulations and real-world experiments. The focus is on BO training efficiency in simulations and the vehicle’s mean projected velocity (Mean($v_{\textrm{p}}$)), lap time (L.T), and computation time (C.T) during real-world racing using a self-built F1TENTH vehicle (see Fig. 4a). A complex racetrack with four sharp corners (as shown in Fig. 4b) is built for simulation and real-world racing. The Cartographer ROS[23] is used for mapping and localization (see Fig. 4c).
V-A Effectiveness of the proposed OFR in simulation
@@
The training efficiency is calculated as $\frac{N_{\textrm{BL}}-N_{\textrm{opt}}}{N_{\textrm{BL}}}$, where $N_{\textrm{opt}}$ is iterations required for BO convergence. The percentage of racing performance improvement is calculated as $\frac{T_{\textrm{lap,BL}}-T_{\textrm{lap,opt}}}{T_{\textrm{lap,BL}}}$, where $T_{\textrm{lap,opt}}$ is the optimal lap time. The lower bound of $\bm{\theta}$ is $[0.1,1,1,1,1,0.1,1,1,0.01]$, and the upper bound is $[1,50,10,10,10,20,50,20,0.4]$. The final $N_{p}$ used to construct the MPC optimization problem is $\left\lfloor{\bm{\theta}_{1}\cdot D_{\textrm{ref}}}\right\rfloor$. The $D_{\textrm{ref}}$ is 62.8m and ${J}_{\textrm{inf}}$ is 17.6. For Ablation1, ${J}_{\textrm{inf}}$ is 35 because the length of racing trajectories are typically shorter than that of the reference line. The results of BO learning based on different objective functions are shown in Fig. 5a and Table II. Compared with baseline, OFR significantly improves the training efficiency of BO by 42.86%. In addition, the OFR converged lap time is also superior to baseline with a reduction of 9.07%. Ablation1 and Ablation2 show lower training efficiency and less reduction in optimal lap time than the proposed OFR. This demonstrates the effectiveness of the trajectory length minimization and lap time reduction components of the proposed OFR. The training efficiency of Ablation3 is comparable to that of the proposed OFR, and it achieves a shorter lap time. However, the slip angle of the optimal trajectory from Ablation3 is significantly larger than that obtained through OFR-based training (as shown in Fig. 5b). This suggests that limiting the maximum orthogonal distance of the racing trajectory can reduce slip angle. This is because OFR penalizes BO for exploring racing trajectories that deviate significantly from the reference line during training.
@@
Methods1 Mean(L.T) Compare2$\downarrow$ Mean($v_{\textrm{p}}$) Compare3$\uparrow$ Max($v$) Laps Std(C.T) Mean(C.T) Std(Cost)    Std($a_{x}$)
@@
  • 1 The offline racing trajectory with globally minimal curvature[13], based on ideal vehicle handling capabilities, is regarded as vehicle limits.
@@
  • 3 Calculated as $\frac{{\textrm{Mean(}v_{\textrm{p},\textrm{lap}}}\textrm{)}}{{\textrm{Mean(}v_{\textrm{p},\textrm{limits}}\textrm{)}}}$. VPMPCC pushes vehicles to limits through the trade-off between velocity prediction and projected velocity maximization.
(a) Racing trajectories corresponding to three different methods.
(b) Longitudinal velocities corresponding to racing trajectories.
@@
Figure 6: A schematic of trajectories and vehicle velocities during a single racing lap. The racing trajectory illustrates that VPMPCC is closest to the centerline at corner ‘4’, thereby ensuring maximum safety while maintaining racing performance. The velocity diagram shows VPMPCC achieves the highest cornering velocities, benefiting from the balance between RVP-based prediction and velocity maximization in optimal local trajectory planning.
V-B Effectiveness of the proposed VPMPCC on scaled vehicle
@@
The MPCC proposed in [16] and the Velocity-Tracking Adaptive Pure Pursuit (VTAPP) controller proposed in [21] are used for comparison. The optimal parameters of MPCC are also trained through OFR-based BO, which is $[13,0,39.0,5.8,0.6,2.8,0.2,0.5,0.02]$. The $v_{\Delta,\textrm{max}}$ is 10m/s. And MPCC uses the racetrack centerline as the reference line, as discussed in [16]. The optimal parameters of VPMPCC trained by BO based on the proposed OFR is $\bm{\theta}^{*}=[6,3,6,3.9,1,19.0,28,15.7,0.3]$. These three methods are used to perform consecutive laps of racing to verify the superiority of the VPMPCC method in terms of high performance and stability. The statistical data are shown in Table III. The Mean($v_{\textrm{p}}$) is calculated as Mean($\frac{D_{\textrm{ref}}}{\textrm{L.T}}$).
@@
The mean projected velocity of VPMPCC reaches 93.18% of vehicle handling limits, much higher than the other methods. It also has a low computation time of 7.04ms, which indicates that VPMPCC has high real-time performance. VPMPCC has consistently raced for 25 laps at a maximum velocity of 6.39 m/s, demonstrating its stability. This successfully proves that VPMPCC pushes the vehicle performance to its limits with stability. Compared to VTAPP, VPMPCC maintains the effect of maximum projected velocity, thus planning local trajectories with superior velocities, especially in corners (as shown in Fig. 6a). As shown in Fig. 6b, during a single lap, the racing velocities of VPMPCC are significantly higher than those of MPCC. This is because MPCC employs an aggressive racing strategy, but the tracking error $\mathbf{e}_{s}$ is not sufficient to enable it to decelerate in time in corner ‘4’, resulting in a possible collision. As a result, the high-velocity racing parameters of MPCC are restricted in BO training, leaving only parameters with limited racing effects. However, VPMPCC can successfully plan optimal local trajectories in corners by incorporating velocity prediction. This reduces cornering time while ensuring collision-free maneuvers, allowing VPMPCC to achieve higher global racing velocities and ultimately reduce lap time. This is also proved by optimal costs of MPCC and VPMPCC in Fig. 7, while MPCC successfully passes corner ‘4’, its parameters cause a premature deceleration in corners ‘1’ to ‘3’ (pointed out by $\uparrow$ in Fig. 6b), leading to a loss in racing performance. In contrast, VPMPCC achieves an optimal velocity profile at each sharp corner. Additionally, the cost changes of VPMPCC are smoother compared to MPCC, which further enhances the efficiency of solving the optimization problem (as shown in Fig. 5c ).
Figure 7: Demonstration of the cost for MPCC and VPMPCC in real vehicle racing. High costs indicate lower velocities. MPCC cannot plan optimal velocity profiles at corners, leading to premature deceleration at corners ‘1’ to ‘3’. In contrast, VPMPCC utilizes velocity prediction to balance projected velocity maximization, maintaining optimal velocities at each corner.
@@ VI CONCLUSIONS
This paper proposes a novel data-driven framework for aggressive autonomous racing that pushes the vehicle to its handling limits. By integrating velocity prediction into local trajectory planning, the vehicle’s mean projected velocity reaches 93.18% of its handling limits. The training efficiency of BO is improved by 42.86% through the proposed objective function adapted to racing. By filtering out unqualified racing trajectories during BO training, the parameters trained in simulation can be directly transferred to a real-world vehicle without retraining. Future work will explore applying VPMPCC to complex multi-vehicle racing scenarios and integrating deep reinforcement learning for decision-making.